### v1

# Discrete Brunn–Minkowski inequality for subsets of the cube

Abstract.
We show that for all subsets $A_{1},A_{2},\dots,A_{n}\subseteq\{0,1\}^{d}$, we have
$|A_{1}+A_{2}+\dots+A_{n}|\geqslant\prod_{i=1}^{n}|A_{i}|^{(\log_{2}(n+1))/n}.$
In addition, we show that for all $A,B\subseteq\{0,1,2\}^{d}$ we have
$|A+B|\geqslant(|A||B|)^{\log(5)/(2\log(3))}.$
We also show that for all finite $A,B\subset\mathbb{Z}^{d}$, and any $V\subseteq\{0,1\}^{d}$ the inequality
$|A+B+V|\geqslant|A|^{1/p}|B|^{1/q}|V|^{\log_{2}(p^{1/p}q^{1/q})}$
holds for all $p\in(1,\infty)$, where $q=\frac{p}{p-1}$ is the conjugate exponent of $p$. All three estimates are dimension free with the best possible exponents. We also discuss applications to various related problems.
2020 Mathematics Subject Classification:
11B30, 11B13,
1. Introduction
1.1. Minkowski sum for subsets of a discrete cube
It is well-known that given arbitrary compact sets $A_{1},\ldots,A_{n}\subset\mathbb{R}^{d}$ the classical Brunn–Minkowski inequality states that
$\mu(A_{1}+\ldots+A_{n})\geqslant\left(\sum_{j=1}^{n}\mu(A_{j})^{1/d}\right)^{d},$ (1.1)
where $d\mu$ is the Lebesgue measure on the product space $\mathbb{R}^{d}$, and $\sum_{j=1}^{n}A_{j}$ denotes the Minkowski sum. The Gaussian counterpart of the inequality (1.1), i.e., the so called Ehrhard’s inequality states that
$\gamma(A_{1}+\ldots+A_{n})\geqslant\Phi\left(\sum_{j=1}^{n}\Phi^{-1}(\gamma(A_{j}))\right),$ (1.2)
[…]
However, less is known when $A_{j}$ are subsets of a discrete cube and the underlying measure is the uniform counting measure. One of the goals of this paper is to find the largest possible power $p_{n,m}$ such that the inequality
$\left|\sum_{j=1}^{n}A_{j}\right|\geqslant\prod_{j=1}^{n}|A_{j}|^{p_{n,m}}$ (1.3)
holds for all subsets $A_{1},\ldots,A_{n}\subset\mathcal{C}_{m}^{d}:=\{0,1,\ldots,m\}^{d}\subset\mathbb{Z}^{d}$, all $d\geqslant 1$, where $n,m$ are fixed nonnegative integers, and ||$\cdot$|| ⋅ | denotes the uniform counting measure of a discrete set.
Let us make a couple of observations.
1. For arbitrary sets one has the following trivial bounds
$\prod_{i=1}^{n}|A_{i}|\geqslant\left|\sum_{j=1}^{n}A_{j}\right|\geqslant\max_{1\leqslant i\leqslant n}\{|A_{i}|\}\geqslant\prod_{i=1}^{n}|A_{i}|^{1/n},$ (1.4)
which imply that $1\geqslant p_{n,m}\geqslant\frac{1}{n}.$
2. Choosing $A_{1}=\ldots=A_{n}=\mathcal{C}_{m}^{d}$ we obtain
$\frac{\log(nm+1)}{n\log(m+1)}\geqslant p_{n,m}\,.$ (1.5)
3. The main result in [17] states that $p_{2,1}=\frac{\log(3)}{\log(4)}$, i.e., we have equality in (1.5) in the case $(n,m)=(2,1)$.
4. For $m\geqslant 3$ let $\tau_{m}$ be the solution of the equation
$1+(m(m+1))^{\tau_{m}}=(m+1)^{2\tau_{m}}.$
The lower bound $p_{2,m}\geqslant\tau_{m}$ was obtained in [1], and it was also asked whether we have equality in (1.5) for $n=2$. The behaviour of the power $\tau_{m}$, as $m$ tends to $\infty$, asymptotically matches the behaviour of the power $\frac{\log(2m+1)}{2\log(m+1)}$.
For any $p\geqslant 1$ and any $f:\mathbb{Z}^{d}\to\mathbb{R}$ with finite support set
$\|f\|_{\ell^{p}(\mathbb{Z}^{d})}=\left(\sum_{x\in\mathbb{Z}^{d}}|f(x)|^{p}\right)^{1/p}.$
In what follows we will write $\|f\|_{p}$ instead of $\|f\|_{\ell^{p}(\mathbb{Z}^{d})}$ and it will be clear from the context what is the value for $d$. Given functions $f,g:\mathbb{Z}^{d}\to\mathbb{R}$ with finite support we define their sup convolution as follows
$f\bar{*}g(z)=\sup_{x\in\mathbb{Z}^{d}}f(x)g(z-x).$
In the next theorem we show that we have equality in (1.5) for the pairs $(n,m)=(n,1)$ for all $n\geqslant 1$.
Theorem 1.1.
For all $d\geqslant 1$, and all $f_{j}:\mathbb{Z}^{d}\to[0,\infty)$, $j=1,\ldots,n$, supported on $\{0,1\}^{d}$, we have
[…]
In particular, for all subsets $A_{j}\subseteq\{0,1\}^{d}$, we have
$\left|\sum_{j=1}^{n}A_{j}\right|\geqslant\left(\prod_{j=1}^{n}|A_{j}|\right)^{p}$ (1.7)
with the best possible power $p=\frac{\log_{2}(n+1)}{n}$.
The inequality (1.7) follows from (1.6) by taking $f_{j}=1_{A_{j}}$ for each $j=1,\ldots,n$. A related inequality to (1.6) was established by Ivanisvili [8, Theorem 1].
Next we affirmatively answer the question asked in [11] which also confirms the case $m=2$ of the question in [1].
Theorem 1.2.
For all $d\geqslant 1$, and all $f,g:\mathbb{Z}^{d}\to[0,\infty)$ supported on $\{0,1,2\}^{d}$, we have
[…]
In particular, for all $A,B\subset\{0,1,2\}^{d}$, and all $d\geqslant 1$ we have
$|A+B|\geqslant|A|^{p}|B|^{p}$
holds with the best possible power $p=\frac{\log(5)}{2\log(3)}=0.732\ldots$.
The following result was obtained by Matolcsi, Ruzsa, Shakan and Zhelezov.
Theorem 1.3 ([13]).
Let $d\geqslant 1$ and $A,B\subset\mathbb{Z}^{d}$ be finite sets and let $V\subseteq\{0,1\}^{d}$. Then
$|A+B+V|\geqslant\left(\frac{p^{1/p}q^{1/q}}{2}\right)^{d}|V||A|^{1/p}|B|^{1/q},$ (1.9)
for any $p\in(1,\infty),$ where $q=\frac{p}{p-1}$ is the conjugate exponent of $p$.
Notice that by choosing $V=\{0,1\}^{d}$ and $1/p=|A|^{1/d}(|A|^{1/d}+|B|^{1/d})^{-1}$ the estimate (1.9) becomes
$|A+B+\{0,1\}^{d}|^{1/d}\geqslant|A|^{1/d}+|B|^{1/d}.$ (1.10)
The inequality (1.10) can be considered as the discrete Brunn–Minkowski inequality as it implies the Brunn–Moinkowski inequality (1.1) for two sets (see [13] and references therein). In this paper we obtain
Theorem 1.4.
For all $d\geqslant 1$, and all finitely supported $f,g,h:\mathbb{Z}^{d}\to[0,\infty)$ with $f$ supported on $\{0,1\}^{d}$ we have
$\|f\bar{*}g\bar{*}h\|_{1}\geqslant\|f\|_{c}\|g\|_{p}\|h\|_{q}$ (1.11)
holds for all $p\in(1,\infty)$, $q=\frac{p}{p-1}$ is the conjugate exponent of $p$, and $c$ satisfies $2^{1/c}=p^{1/p}q^{1/q}$. In particular, we have
$|V+A+B|\geqslant|V|^{1/c}|A|^{1/p}|B|^{1/q},$ (1.12)
for all finite subsets $A,B\subset\mathbb{Z}^{d}$ and $V\subset\{0,1\}^{d}$.
The theorem positively answers the question asked in [15] (see the discussion right after Theorem 11.1 in [13], and see [15]). Notice that Theorem 1.4 implies Theorem 1.3 because of the trivial estimate
$|V|^{1/c}\geqslant\left(\frac{p^{1/p}q^{1/q}}{2}\right)^{d}|V|$
due to $c\geqslant 1$ and $|V|^{1-1/c}\leqslant|\{0,1\}^{d}|^{1-1/c}=\left(\frac{2}{p^{1/p}q^{1/q}}\right)^{d}$.
Theorem 1.3, in the case $p=q=2$, is the key ingredient in obtaining query complexity for learning subsets $A\subset\mathbb{Z}^{d}$ having small doubling constant, see [18]. A short self-contained proof of Theorem 1.3 was recently given in [5] in the case $p=q=2$.
For any positive integer $k$ set
$kA:=\underbrace{A+A+\dots+A}_{k}.$
Instead of considering subsets of cubes, one may consider sets containing cubes, that is the content of our next result.
Theorem 1.5.
Let $d\geqslant 1$. The following inequalities hold
$\left(k-1+\frac{1}{k}\right)^{d}\geqslant\inf_{\{0,1,\dots,k-1\}^{d}\subseteq A\subset\mathbb{Z}^{d}}\frac{|kA|}{|A|}\geqslant k^{\frac{d(k-1)}{k}}.$
Remark 1.
Observe that $\lim_{k\to\infty}\frac{(k-1+\frac{1}{k})^{d}}{k^{\frac{d(k-1)}{k}}}=1$, thus our bounds are asymptotically optimal.
Remark 2.
Our Theorem 1.5 can be seen as an extension of a result obtained by Green and Tao [6, Proposition 4.1], for $k=2$.
For any $\lambda>0$ set
$A+\lambda\cdot B=\{a+\lambda b;\ a\in A,b\in B\}.$
The additive energy $E(A)$ of a finite subset $A$ of an additive group $G$ is defined as $E(A):=|\{(a_{1},a_{2},a_{3},a_{4})\in A^{4};\ a_{1}+a_{2}=a_{3}+a_{4}\}|,$ (see [16]). It was proved by Kane and Tao [10, Theorem 7] that $E(A)\leqslant|A|^{\log_{2}6}$ for all $A\subseteq\{0,1\}^{d}$. For any $k\geqslant 2$, the $k-$additive energy $E_{k}(A)$ of a finite subset $A$ of an additive group $G$ is defined as $E_{k}(A):=|\{(a_{1},a_{2},\dots,a_{2k})\in A^{2k};\ a_{1}+a_{2}+\dots+a_{k}=a_{k+1}+a_{k+2}+\dots+a_{2k}\}|.$ The result of Kane-Tao was later extended to all $k\geqslant 2$, see [2, Theorem 2] and [12, Corollary 2], establishing that $E_{k}(A)\leqslant|A|^{\log_{2}{{2k}\choose{k}}}$ for all $A\subseteq\{0,1\}^{d}$. Combining this estimate with our Theorem 1.1 we obtain
$E_{k}(A)$ $\leqslant|A|^{\log_{2}{{2k}\choose{k}}}=|A|^{(\log_{2}{\frac{k+1}{2}})(\log_{\frac{k+1}{2}}{{2k}\choose{k}})}$
[…]
The upper bound on $E_{k}(A)$ in terms of $|kA|$ gives a relation between additive energies and doubling constants.
Corollary 1.
Let $d\geqslant 1$, $k\geqslant 2$, and $p=\log_{\frac{k+1}{2}}{{2k}\choose{k}}$. We have
$E_{k}(A)\leqslant\left(\frac{|kA|}{|A|}\right)^{p}$
for all $A\subseteq\{0,1\}^{d}$. Moreover, $p$ can not be replaced by a smaller quantity. In particular, $E(A)\leqslant\left(|A+A|/|A|\right)^{\log_{3/2}6}$ for all $A\subseteq\{0,1\}^{d}$.
The sharpness of the power $p$ in the corollary follows by testing the inequality for $A=\{0,1\}^{d}$.
In [7] Hanson and Petridis study sum of dilates, they proved that if $A$ is a subset of a commutative group $G$ and $K$ is a parameter such that $|A+A|\leqslant K|A|$, then $|A+2\cdot A|\leqslant K^{2.95}|A|$.
As a consequence of the simple inequality $\frac{|A+2\cdot A|}{|A|}\leqslant|A|$ and our Theorem 1.1 we obtain
Corollary 2.
Let $d\geqslant 1$ and $p=\log_{3/2}2\approx 1.7095$. The inequality
$|A+2\cdot A|\leqslant\left(\frac{|A+A|}{|A|}\right)^{p}|A|$
holds for all $A\subseteq\{0,1\}^{d}$. Moreover, $p$ cannot be replaced by a smaller power.
The sharpness of the power $p=\log_{3/2}2$ follows by considering $A=\{0,1\}^{d}$.
Next we study a variant of the distinct distances problem posed by Erdős. For positive integers $N,k,\ell,m$, we consider sets of $N$ points in $\mathbb{R}^{2}$ with the property that every $k$ of the points span at least $\ell$ distinct $m$-sums. Let $\psi_{m}(N,k,\ell)$ denote the minimum number of $m-$sums such a point set can span.
Corollary 3.
For every $N\geqslant k>0$ and $m\geqslant 1$, where $N$ is the power of two, we have
$\psi_{m}\left(N,k,k^{\log_{2}(m+1)}\right)=O\left(N^{\log_{2}(m+1)}\right).$
This can be seen as an extension of the main theorem in [3]. As in [3] the hypercube $\{0,1\}^{n}\subset\mathbb{R}^{n}$ can be projected (which is also bijection onto its image) by a transformation $T$ onto the real line $\mathbb{R}$ giving a set $P_{n}$ of cardinality $N=2^{n}$ so that any subset of $\tilde{A}\subset P_{n}$ has the same number of distinct m-sums as the set $A\subset\{0,1\}^{n}$, where $T(A)=\tilde{A}$. Thus the corollary is a direct consequence of Theorem 1.1. The projection is defined as follows: Let $n\in\mathbb{Z}^{+}$. Choose $r_{1},r_{2},\dots,r_{n}\in\mathbb{R}\setminus\mathbb{Q}$ such that $r_{i}>0$ for all $1\leqslant i\leqslant n$ and the only integer solution to $x_{1}\cdot r_{1}+x_{2}\cdot r_{2}+\cdots+x_{n}\cdot r_{n}=0,$ is $x_{1}=x_{2}=\cdots=x_{n}=0$. For every integer $0\leq j\leq n$, we define a set $P_{j}$ of $2^{j}$ real numbers, as follows. We set $P_{0}=\{1\}$. For $1\leq j\leq n$, we define $P_{j}=P_{j-1}\bigcup(r_{j}+P_{j-1})$.
In the next section we present the proofs of the main results.
2. Proof of the main results
The next theorem shows the functional inequalities (1.6), (1.8) and (1.11) hold for all dimensions $d\geqslant 1$ if and only if they hold in dimension $d=1$.
Theorem 2.1 (Compressing dimension).
Fix positive numbers $p_{1},\ldots,p_{n}>0$ and nonempty finite sets $S_{1},\ldots,S_{n}\subset\mathbb{Z}$. The inequality
$\sum_{z\in\mathbb{Z}^{d}}\max_{x^{(1)}+\ldots+x^{(n)}=z}\prod_{j=1}^{n}f^{p_{j}}_{j}(x^{(j)})\geqslant\prod_{j=1}^{n}\left(\sum_{x\in\mathbb{Z}^{d}}f_{j}(x)\right)^{p_{j}}$ (2.1)
holds for all non-negative functions $f_{j}$ supported on $S_{j}^{d}$, $j=1,\ldots,n$, and all $d\geqslant 1$, if and only if it holds for $d=1$.
Proof.
Suppose (2.1) holds for $d=1$. For each $j=1,\ldots,n,$ let $x^{(j)}=(\overline{x}^{(j)},\tilde{x}^{(j)})\in\mathbb{Z}^{d}$, where $\overline{x}^{(j)}\in\mathbb{Z}^{d-1}$. Similarly let $z=(\overline{z},\tilde{z})\in\mathbb{Z}^{d}$. Set $\overline{f_{j}}(\overline{x}^{(j)})=\sum_{\tilde{x}^{(j)}\in\mathbb{Z}}f_{j}(x^{(j)})$ for each $j=1,\ldots,n$. We have
$\sum_{z\in\mathbb{Z}^{d}}\max_{x^{(1)}+\ldots+x^{(n)}=z}\prod_{j=1}^{n}f^{p_{j}}_{j}(x^{(j)})=$
$\sum_{\overline{z}\in\mathbb{Z}^{d-1}}\sum_{\tilde{z}\in\mathbb{Z}}\max_{\overline{x}^{(1)}+\ldots+\overline{x}^{(n)}=\overline{z}}\max_{\tilde{x}^{(1)}+\ldots+\tilde{x}^{(n)}=\tilde{z}}\prod_{j=1}^{n}f^{p_{j}}_{j}(\overline{x}^{(j)},\tilde{x}^{(j)})\geqslant$
$\sum_{\overline{z}\in\mathbb{Z}^{d-1}}\max_{\overline{x}^{(1)}+\ldots+\overline{x}^{(n)}=\overline{z}}\sum_{\tilde{z}\in\mathbb{Z}}\max_{\tilde{x}^{(1)}+\ldots+\tilde{x}^{(n)}=\tilde{z}}\prod_{j=1}^{n}f^{p_{j}}_{j}(\overline{x}^{(j)},\tilde{x}^{(j)})\stackrel{{\scriptstyle\text{case}\,d=1}}{{\geqslant}}$
$\sum_{\overline{z}\in\mathbb{Z}^{d-1}}\max_{\overline{x}^{(1)}+\ldots+\overline{x}^{(n)}=\overline{z}}\prod_{j=1}^{n}\overline{f}^{p_{j}}_{j}(\overline{x}^{(j)})\stackrel{{\scriptstyle\text{iterate}}}{{\geqslant}}\ldots\geqslant\prod_{j=1}^{n}\left(\sum_{x\in\mathbb{Z}^{d}}f_{j}(x)\right)^{p_{j}}.$
In the case $d=1$, $S_{1}=\ldots=S_{n}=\{0,1,\ldots,m\}$, after a change of variables, the inequality (2.1) can be rewritten as follows: for any non-negative $y_{i}^{(j)}$ with $i=0,\ldots,m$, and $j=1,\ldots,n,$ we have
$\sum_{k=0}^{nm}\max_{i_{1}+\dots+i_{n}=k}y_{i_{1}}^{(1)}y_{i_{2}}^{(2)}\dots y_{i_{n}}^{(n)}\geqslant\prod_{j=1}^{n}\left(\sum_{i=0}^{m}(y_{i}^{(j)})^{1/p_{j}}\right)^{p_{j}}.$ (2.2)
It turns out that to verify (2.2) for some $p_{1},\ldots,p_{n}>0$ it suffices to consider the case when for each $j=1,\dots,n,$ the sequence $\{y_{i}^{(j)}\}_{i=0}^{m}$ is non-increasing.
Lemma 2.2.
Let $y_{i}^{(j)}$ with $i=0,1,\dots,m$ and $j=1,2,\dots,n$ be non-negative real numbers. Then if for each $j=1,\dots,n$ we sort all $x_{i}^{(j)}$ from the largest to the smallest, the value of the left hand side of (2.2) will not increase.
Proof.
It suffices  [^1] Indeed, this follows from the fact that the sum of numbers can be written as an integral of the cardinality of the set of indices corresponding to numbers at least $t$. to show that for any $t\geqslant 0$ the cardinality of the set
$\mathcal{K}(t):=\{k:\,\exists i_{1}+\dots+i_{n}=k,\,x_{i_{1}}^{(1)}x_{i_{2}}^{(2)}\dots x_{i_{n}}^{(n)}\geqslant t\}\subset\{0,1,\dots,nm+1\}$
will not increase. This can be written as an infinite union
$\cup_{t_{1}\dots t_{n}=t}\{i_{1}:x_{i_{1}}^{(1)}\geqslant t_{1}\}+\dots+\{i_{n}:x_{i_{n}}^{(n)}\geqslant t_{n}\},$
where the sum is understood as a sum of subsets of $\mathbb{Z}$. Rearranging $x_{i}^{(j)}$’s for each $j$ does not change the cardinality of the sets involved. After the sort each of the sets becomes an interval of the form $[0,L)$ and so the union of all these sums is equal to its maximal value which, in turn, is equal to
$\max_{t_{1}\dots t_{n}=t}1-n+\sum_{j=1}^{n}\left|\left\{x_{i_{j}}^{(j)}\geqslant t_{j}\right\}\right|,$
which is a lower bound for the original value as follows from the Cauchy-Davenport theorem. ∎
Remark 3.
To prove Theorem 1.1, Theorem 1.2, and Theorem 1.4 it suffices to verify (2.1) in the case $d=1$, A) $S_{1}=S_{2}=\ldots=S_{n}=\{0,1\}$, $p_{1}=\ldots=p_{n}=\frac{\log_{2}(n+1)}{n}$; B) $n=2$, $S_{1}=S_{2}=\{0,1,2\}$, $p_{1}=p_{2}=\frac{\log(5)}{2\log(3)}$; C) $n=3$, $S_{1}=\{0,1\}$, $S_{2},S_{3}\subset\mathbb{Z}$ are arbitrary finite, $p_{1}=1/c,p_{2}=1/p,p_{3}=1/q$.
2.1. The proof of Theorem 1.1
Our goal is to prove the following inequality
$\sum_{k=0}^{n}\left(\max_{i_{1}+\dots+i_{n}=k}y_{i_{1}}^{(1)}y_{i_{2}}^{(2)}\dots y_{i_{n}}^{(n)}\right)^{p}\geqslant\prod_{j=1}^{n}\left(y_{0}^{(j)}+y_{1}^{(j)}\right)^{p}.$ (2.3)
By homogeneity we can assume that for each $j=1,\ldots,n$ we have $y_{0}^{(j)}+y_{1}^{(j)}=1$. Also without loss of generality we can assume $y_{0}^{(j)}\geqslant y_{0}^{(j+1)}$ for $j=1,\ldots,n-1$, and $y_{0}^{(1)}\geqslant 1/2$. It will be convenient to denote $x_{i,0}=y^{(i)}_{0}$ for all $i=1,\ldots,n$, and also set $q=pn$. The desired inequality takes the form
$\prod_{i=1}^{n}x^{q/n}_{i,0}+\prod_{i=1}^{n}(1-x_{i,0})^{q/n}+\sum_{\ell=1}^{n-1}\prod_{i=1}^{\ell}x^{q/n}_{i,0}\prod_{i=\ell+1}^{n}(1-x_{i,0})^{q/n}\geqslant 1.$ (2.4)
For each $i=1,\ldots,n$, we define
$w_{i}=\left[\frac{1-x_{n+1-i,0}}{x_{n+1-i,0}}\right]^{q/n}.$
The inequality (2.4) reduces to verify that
$1+w_{1}+w_{1}w_{2}+w_{1}w_{2}w_{3}+\dots+\prod_{i=1}^{n}w_{i}\geqslant\prod_{i=1}^{n}(1+w^{n/q}_{i})^{q/n}.$ (2.5)
[…]
Lemma 2.3.
We have
$\left(1+w_{1}+w_{1}w_{2}+w_{1}w_{2}w_{3}+\dots+\prod_{i=1}^{n}w_{i}\right)^{n}\geqslant\prod_{i=1}^{n}\left(\sum_{k=0}^{n}w^{k}_{i}\right).$
for any nonnegative numbers $w_{1},\ldots,w_{n},$ with $w_{1}\geqslant w_{2}\geqslant\ldots\geqslant w_{n}$.
For each positive integer $k$ consider $f_{k}(w)=(1+w+\dots+w^{k})^{1/k}$ on $\mathbb{R}_{+}$. We need the following technical statement.
Lemma 2.4.
For every $n\geqslant 2$ let $g_{n}(w)=f_{n-1}(w)/f_{n}(w)$. Then $g_{n}$ satisfies $g_{n}(1/w)=g_{n}(w)$, is increasing on $[0,1]$, and decreasing on $[1,\infty)$.
Proof of Lemma 2.4.
The equality $g_{n}(w)=g_{n}(1/w)$ follows from the fact that $f_{k}(1/w)=f_{k}(w)/w$ for positive integers $k$. By the $w\rightarrow 1/w$ symmetry it suffices to prove that the function is decreasing for $w\geqslant 1$. Taking logarithmic derivative we observe that $(\ln(g_{n}(w)))^{\prime}\leqslant 0$ is equivalent to have the inequality
$\frac{1}{n}\cdot\frac{1+2w+\dots+nw^{n-1}}{1+w+\dots+w^{n}}\geqslant\frac{1}{n-1}\cdot\frac{1+2w+\dots+(n-1)w^{n-2}}{1+w+\dots+w^{n-1}}.$
After multiplication by denominators and substraction of the common part, i.e. subtracting $(n-1)(1+2w+\dots+(n-1)w^{n-2})(1+w+\dots+w^{n-1})$, we get
$(n-1)nw^{n-1}(1+w+\dots+w^{n-1})$
$\geqslant(1+w+\dots+w^{n-1})(1+2w+\dots+(n-1)w^{n-2})$
$\ \ \ \ +nw^{n}(1+2w+\dots+(n-1)w^{n-2}).$
Expanding both sides and subtracting we arrive at the inequality
$\sum_{k=1}^{n-1}\frac{k(k+1)}{2}\cdot w^{2n-2-k}\geqslant\sum_{k=1}^{n-1}\frac{k(k+1)}{2}\cdot w^{k-1},$
which is trivial for $w\geqslant 1$. ∎
Proof of Lemma 2.3.
We take the $n$-th root of both sides, so the inequality gets rewritten as
$1+w_{1}+w_{1}w_{2}+\dots+w_{1}w_{2}\dots w_{n}\geqslant\prod_{i=1}^{n}f_{n}(w_{i}).$
We will use induction on $n$, the case $n=1$ is trivial. Take $n\geqslant 2$ and assume the inequality is known for $n-1$ variables. Note that the inequality is invariant under the change $(w_{1},\dots,w_{n})\rightarrow(1/w_{n},\dots,1/w_{1})$. So making this change, if needed, we may assume that $w_{1}w_{n}\geqslant 1$, which we do from now on.
First, we apply the inequality to $(w_{2},\dots,w_{n})$ and write $w$ in place of $w_{1}$, so that the inequality is reduced to the following:
$1+w\cdot\prod_{i=2}^{n}f_{n-1}(w_{i})\geqslant f_{n}(w)\cdot\prod_{i=2}^{n}f_{n}(w_{i}).$
Now, we write $w$ as $(f_{n}(w)^{n}-1)/f_{n-1}(w)^{n-1}$ and divide both sides by $f_{n}(w)^{n}$ to arrive at an equivalent inequality
$\frac{1}{f_{n}(w)^{n}}+\left(1-\frac{1}{f_{n}(w)^{n}}\right)\cdot\prod_{i=2}^{n}\frac{f_{n-1}(w_{i})}{f_{n-1}(w)}\geqslant\prod_{i=2}^{n}\frac{f_{n}(w_{i})}{f_{n}(w)}.$
Note that since both $f_{n-1}$ and $f_{n}$ are increasing and $w\geqslant w_{i}$, both products are at most $1$, and since the left hand side of the inequality above is a convex combination of the product and 1, it suffices to prove that
$\prod_{i=2}^{n}\frac{f_{n-1}(w_{i})}{f_{n-1}(w)}\geqslant\prod_{i=2}^{n}\frac{f_{n}(w_{i})}{f_{n}(w)},$
which would follow from the inequality term-wise, which, in turn, can be rewritten as
$\frac{f_{n-1}(w_{i})}{f_{n}(w_{i})}\geqslant\frac{f_{n-1}(w)}{f_{n}(w)}.$
This is true since we have $ww_{n}\geqslant 1$, which implies that $w\geqslant 1/w_{n}\geqslant 1/w_{i}$ and so taking $y:=\max\{w_{i},1/w_{i}\}$ we have $1\leqslant y\leqslant w$ and by Lemma 2.4
$\frac{f_{n-1}(w_{i})}{f_{n}(w_{i})}=\frac{f_{n-1}(y)}{f_{n}(y)}\geqslant\frac{f_{n-1}(w)}{f_{n}(w)}.$
Lemma 2.5.
Let $r=\log_{2}(n+1)$. We have
$\sum_{i=0}^{n}w^{i}-(1+w^{n/r})^{r}\geqslant 0$
for all $w\geqslant 0$.
We will make a general observation which implies Lemma 2.5, and which also will be used later in proofs of other technical statements. The general observation says that if we consider the sequence of polynomials
$P_{n}(x)=1+x+x^{2}+\ldots+x^{n},$
then the functions $(P_{n}(x))^{1/n}$ and $(P_{m}(x))^{1/m}$ behave similarly at infinity. We can match their values at $x=1$ by replacing $(P_{m}(x))^{1/m}$ with $P_{m}(x^{s})^{1/(ms)},$ where $s=\frac{n\log(m+1)}{m\log(n+1)}$ solves the equation $(P_{n}(1))^{1/n}=P_{m}(1)^{1/(ms)}$. Then it turns out that $(P_{n}(x))^{1/n}\geqslant P_{m}(x^{s})^{1/(ms)}$ for all $x\geqslant 0$ if and only if $n\geqslant m$. The inequality $(P_{n}(x))^{1/n}\geqslant(P_{1}(x^{s}))^{1/s}$ is the same as in Lemma 2.5. The general observation follows from the following
Lemma 2.6.
For any $n>m\geqslant 1$ and $x>0$ one has $F_{n}(x)\geqslant F_{m}(x)$, where
$F_{n}(x):=P_{n}\left(x^{\tfrac{\log{(n+1)}}{n}}\right)^{\tfrac{1}{\log{(n+1)}}}.$
In order to prove the lemma we use the following
Proposition 2.7.
For every $n\geqslant 1$, the function $F_{n}$ satisfies $F_{n}(0)=1,F_{n}(1)=e,F_{n}(1/x)=F_{n}(x)/x,F_{n}^{\prime}(1)=e/2$.
Proof.
First three assertions are obvious and the last one follows by differentiating the functional equation and substituting $x=1$. ∎
Proof of Lemma 2.6.
Without loss of generality we may assume that $n=m+1$. By Proposition 2.7 the equation $F_{n}(x)=F_{m}(x)$ has a root at $x=0$, a double root at $x=1$ and all other roots come in pairs $(x,1/x)$. Since close to zero we have $F_{n}(x)\sim 1+x^{\tfrac{\log{(n+1)}}{n}}/\log{(n+1)}$, we know that $F_{n}(x)\geqslant F_{m}(x)$ in some neighbourhood of zero. Hence, if the statement of the lemma is false, we must have at least one $t\in(0,1)$ for which $F_{n}(t)=F_{m}(t)$. Seeking a contradiction we assume that $F_{n}(t)=F_{m}(t)$ for some $t\in(0,1)$.
Let us compute the logarithmic derivative of $F_{n}$.
$\partial_{x}\log F_{n}(x)$ $=\frac{1}{\log{(n+1)}}\times$
$\left(-\frac{\tfrac{(n+1)\log{(n+1)}}{n}\cdot x^{\tfrac{(n+1)\log{(n+1)}}{n}-1}}{1-x^{\tfrac{(n+1)\log{(n+1)}}{n}}}+\frac{\tfrac{\log{(n+1)}}{n}\cdot x^{\tfrac{\log{(n+1)}}{n}-1}}{1-x^{\tfrac{\log{(n+1)}}{n}}}\right)$
$=\frac{1}{nx}\cdot\left(-\frac{(n+1)\cdot x^{\tfrac{(n+1)\log{(n+1)}}{n}}}{1-x^{\tfrac{(n+1)\log{(n+1)}}{n}}}+\frac{x^{\tfrac{\log{(n+1)}}{n}}}{1-x^{\tfrac{\log{(n+1)}}{n}}}\right)$
$=\frac{1}{nx}\cdot\left(-(n+1)\cdot\left[\frac{1}{1-x^{\tfrac{(n+1)\log{(n+1)}}{n}}}-1\right]\right.$
$+\left.\left[\frac{1}{1-x^{\tfrac{\log{(n+1)}}{n}}}-1\right]\right)$
$=\frac{1}{x}+\frac{1}{x}\cdot\left(-\frac{(n+1)/n}{1-x^{\tfrac{(n+1)\log{(n+1)}}{n}}}+\frac{1/n}{1-x^{\tfrac{\log{(n+1)}}{n}}}\right)$
$=\frac{1}{x}+\frac{1}{x}\cdot\frac{-1+\tfrac{n+1}{n}\cdot x^{\beta_{n}}-\tfrac{1}{n}\cdot x^{\alpha_{n}}}{(1-x^{\alpha_{n}})(1-x^{\beta_{n}})},$
where
$\alpha_{n}:=\tfrac{(n+1)\log{(n+1)}}{n},\quad\text{and \quad}\beta_{n}:=\tfrac{\log{(n+1)}}{n}.$
We now consider the function
$H_{n,m}(x):=$
$x\cdot(\partial_{x}\log F_{n}(x)-\partial_{x}\log F_{m}(x))\cdot(1-x^{\alpha_{n}})(1-x^{\beta_{n}})(1-x^{\alpha_{m}})(1-x^{\beta_{m}}),$
which can be written as a pseudo-polynomial
$H_{n,m}(x)=(1-x^{\alpha_{m}})(1-x^{\beta_{m}})\left(-1+\tfrac{n+1}{n}\cdot x^{\beta_{n}}-\tfrac{1}{n}\cdot x^{\alpha_{n}}\right)-$
$(1-x^{\alpha_{n}})(1-x^{\beta_{n}})\left(-1+\tfrac{m+1}{m}\cdot x^{\beta_{m}}-\tfrac{1}{m}\cdot x^{\alpha_{m}}\right).$
Expanding $H_{n,m}(x)$ as a pseudo-polynomial in the increasing order of exponents, we get
$H_{n,m}(x)$ $=\tfrac{1}{n}x^{\beta_{n}}-\tfrac{1}{m}x^{\beta_{m}}+\left(\tfrac{m+1}{m}-\tfrac{n+1}{n}\right)x^{\beta_{m}+\beta_{n}}+\tfrac{m+1}{m}x^{\alpha_{m}}-\tfrac{n+1}{n}x^{\alpha_{n}}$
$-\left(\tfrac{n+1}{n}+\tfrac{1}{m}\right)x^{\beta_{n}+\alpha_{m}}-x^{\alpha_{m}+\beta_{m}}+x^{\alpha_{n}+\beta_{n}}+\left(\tfrac{1}{n}+\tfrac{m+1}{m}\right)x^{\beta_{m}+\alpha_{n}}$
$+\tfrac{n+1}{n}x^{\alpha_{m}+\beta_{m}+\beta_{n}}-\tfrac{m+1}{m}x^{\alpha_{n}+\beta_{n}+\beta_{m}}-\left(\tfrac{1}{m}-\tfrac{1}{n}\right)x^{\alpha_{m}+\alpha_{n}}$
$+\tfrac{1}{m}x^{\alpha_{n}+\beta_{n}+\alpha_{m}}-\tfrac{1}{n}x^{\alpha_{m}+\beta_{m}+\alpha_{n}}.$
To see that exponents are increasing, note that all exponents split into pairs with sum $\alpha_{m}+\beta_{m}+\alpha_{n}+\beta_{n}$, so it suffices to show that
$\beta_{n}<\beta_{m}<\beta_{m}+\beta_{n}<\alpha_{m}<\alpha_{n}<\beta_{n}+\alpha_{m}<\alpha_{m}+\beta_{m}<\alpha_{n}+\beta_{n}.$
Inequalities
$\beta_{n}<\beta_{m}<\beta_{m}+\beta_{n}<\alpha_{m}<\beta_{n}+\alpha_{m}<\alpha_{m}+\beta_{m}<\alpha_{n}+\beta_{n},$
are satisfied for any $n>m$, as is the trivial inequality $\alpha_{n}>\alpha_{m}$. To verify that $\alpha_{n}<\beta_{n}+\alpha_{m}$ we recall our assumption that $n=m+1$ and so
$\alpha_{m}=\tfrac{(m+1)\log{(m+1)}}{m}>\log{(m+2)}=\log{(n+1)}=\alpha_{n}-\beta_{n},$
which implies the desired inequality. Here, the inequality in the middle is equivalent to $(m+2)^{m}<(m+1)^{m+1}$ which is true for $m\geqslant 1$.
So the signs of the monomials are ++++++++-++---$+$--+-+ - + + - - - + + + - - + -, making it 7 sign changes. Hence, by the Descartes’ rule of signs for pseudo-polynomials (see example #77 on page 46 in [14]) we conclude that $H_{n,m}$ has at most 7 positive roots counted with multiplicities. Since the product $(1-x^{\alpha_{n}})(1-x^{\beta_{n}})(1-x^{\alpha_{m}})(1-x^{\beta_{m}})$ already has root 1 of multiplicity 4, we conclude that
$\partial_{x}\log F_{n}(x)-\partial_{x}\log F_{m}(x)=\partial_{x}(\log F_{n}(x)-\log F_{m}(x))$
must have at most 3 positive roots counted with multiplicities.
Since $F_{n}(0)=F_{m}(0)$ and $F_{n}(1)=F_{m}(1)$, and, by our assumption, $F_{n}(t)=F_{m}(t)$ for some $t\in(0,1)$, we conclude by Rolle’s theorem that $\partial_{x}(\log F_{n}(x)-\log F_{m}(x))$ has a root in each of $(0,t)$ and $(t,1)$. By the functional equation $F_{n}(1/x)=F_{n}(x)/x$ we have $F_{n}(1/t)=F_{m}(1/t)$ and so $\partial_{x}(\log F_{n}(x)-\log F_{m}(x))$ must also have a root in $(1,1/t)$, again by Rolle’s theorem. Hence, together with $x=1$ the function $\partial_{x}(\log F_{n}(x)-\log F_{m}(x))$ has at least 4 positive roots leading to a contradiction. ∎
2.2. The proof of Theorem 1.2
Using Remark 3 it suffices to verify the following six-point inequality.
Proposition 2.8.
For all $x_{0},x_{1},x_{2}\geq 0$ and $y_{0},y_{1},y_{2}\geq 0$, it holds that
$\sum_{k=0}^{4}\max_{i+j=k}x_{i}^{p}y_{j}^{p}\geq(x_{0}+x_{1}+x_{2})^{p}(y_{0}+y_{1}+y_{2})^{p}\,.$ (2.6)
We now prove Proposition 2.8. We proceed by successively reducing the number of variables, deducing the proposition from the following two lemmas.
Lemma 2.9.
The following inequalities hold for all $x\geq 0$
$1+x^{p}\geq(1+x)^{p}$ (2.7)
$1+x^{p}+x^{2p}\geq(1+x)^{2p}$ (2.8)
$1+x^{p}+x^{2p}+x^{3p}+x^{4p}\geq(1+x+x^{2})^{2p}$ (2.9)
$1+x^{p}+x^{2p}+x^{3p}\geq(1+x)^{p}(1+x+x^{2})^{p}\,.$ (2.10)
Proof.
Inequality (2.7) holds since $p<1$. Inequalities (2.8) and (2.9) follow from Lemma 2.5. We now prove the inequality (2.10). Near 00, the sharper inequality
$1+x^{p}\geq(1+x)^{p}(1+x+x^{2})^{p}$ (2.11)
holds. Indeed, taking logarithmic derivatives of both sides and taking the difference yields
$\frac{1}{p}\partial_{x}\log\left(\frac{1+x^{p}}{(1+x)^{p}(1+x+x^{2})^{p}}\right)=\frac{x^{p-1}}{1+x^{p}}-\frac{1}{1+x}-\frac{1+2x}{1+x+x^{2}}\,.$
The sign of the logarithmic derivative coincides (after clearing all the denominators) with the sign of
$x^{p-1}-2-4x-2x^{1+p}-3x^{2}-2x^{2+p}\,.$
This clearly changes sign exactly once, from positive to negative, and hence (2.11) holds on an interval if it holds at the endpoints. By verifying the inequality at $x=0.1$ we find that it holds in particular on $[0,0.1]$.
It remains to show (2.10) on $[0.1,1]$. From Lemma 2.5, we have that
$1+x^{p}+x^{2p}+x^{3p}\geq(1+x^{3p/2})^{2}\,,$
so it is enough to show that for $x\in[0.1,1]$
$(1+x^{3p/2})^{2}\geq(1+x)^{p}(1+x+x^{2})^{p}\,.$ (2.12)
Again, we compute logarithmic derivatives to obtain
$\frac{1}{p}\partial_{x}\log\left(\frac{(1+x^{3p/2})^{2}}{(1+x)^{p}(1+x+x^{2})^{p}}\right)=3\frac{x^{3p/2-1}}{1+x^{3p/2}}-\frac{1}{1+x}-\frac{1+2x}{1+x+x^{2}}\,.$
After clearing denominators, we see that the sign of the logarithmic derivative coincides with the sign of
$-2+3x^{3p/2-1}-4x+4x^{3p/2}-3x^{2}+2x^{3p/2+1}=:P(x)\,.$ (2.13)
$P$ has at most $5$ zeros on $(0,\infty)$, since this is the number of sign changes in the coefficients. Furthermore $P(x)=-x^{3p/2+1}P(1/x)$, so $P(1)=0$ and the number of zeros of $P$ in $(0,1)$ is the same as the number of zeros in $(1,\infty)$. Clearly $P$ is negative near 00, and one computes that $P^{\prime}(1)<0$. Thus, if there is more than one zero in $(0,1)$, then there are at least $3$, so at least $7$ in total, which is impossible. So inequality (2.12) holds if it holds in $0.1$ and in $1$, which it does. ∎
Lemma 2.10.
The following three inequalities hold for all $0\leq x\leq y$:
$1+y^{p}+y^{2p}+x^{p}y^{2p}+x^{2p}y^{2p}\geq(1+y+xy)^{2p}$ (2.14)
$1+x^{p}+x^{p}y^{p}+x^{2p}y^{p}+x^{2p}y^{2p}\geq(1+x+xy)^{2p}$ (2.15)
$1+y^{p}+x^{p}y^{p}+x^{p}y^{2p}+x^{2p}y^{2p}\geq(1+x+xy)^{p}(1+y+xy)^{p}$ (2.16)
Proof.
Proof of inequality (2.14): We note that the function $f(x):=x^{p}-x^{p-1}-2x^{2p-1}$ is increasing on $[0,\infty)$: Its derivative
$\partial_{x}f(x)=px^{p-1}+(1-p)x^{p-2}-2(2p-1)x^{2p-2}$
is positive by the weighted AM-GM inequality
$px+(1-p)\geq x^{p}>2(2p-1)x^{p}\,,$
where we use that $41/2<p<3/$ 4.
Now we consider $h=1+y+xy$ to be fixed, and show that under this constraint the left hand side of (2.14) is minimized at $x=y$ or $x=0$. To show this, it suffices to prove that its derivative with respect to $y$ changes sign at most once from plus to minus. At the two endpoints $x=y$ and $x=0$ the inequality holds by Lemma 2.9, so this proves the inequality.
We have $\partial_{y}x=-\frac{1+x}{y}$, from which we obtain:
$\frac{y^{1-2p}}{p}\partial_{y}(1+y^{p}+y^{2p}+x^{p}y^{2p}+x^{2p}y^{2p})=y^{-p}+2+x^{p}-x^{p-1}-2x^{2p-1}\,.$ (2.17)
Thus the derivative has the same sign as $y^{-p}+2+f(x)=y^{-p}+2+f(\frac{h-1}{y+1})$, which is decreasing in $y$.
Proof of inequality (2.15): We set $z=xy$. Then we have to show that for $0\leq x^{2}\leq z$, we have
$1+z^{p}+z^{2p}+(1+z^{p})x^{p}\geq(1+x+z)^{2p}\,.$
At the endpoints $x=0$ and $x^{2}=z$, this holds by Lemma 2.9. Near 00, the difference of left hand side and right hand side has derivative $\sim p(1+z^{p})x^{p-1}$, which is positive. So it is enough to show that the logarithmic derivative of the ratio of left hand side and right hand side vanishes at most once. It is
$\frac{p(1+z^{p})x^{p-1}}{1+z^{p}+z^{2p}+(1+z^{p})x^{p}}-2p\frac{1}{1+x+z}\,.$
Multiplying through, it suffices to show that
$\quad p(1+z^{p})x^{p-1}(1+z+x)-2p(1+z^{p}+z^{2p}+(1+z^{p})x^{p})$
$=x^{p-1}p(1+z^{p})(1+z))-2p(1+z^{p}+z^{2p})-x^{p}p(1+z^{p})$
has at most one zero. This follows from Descarte’s rule of signs for pseudo-polynomials, since there is only one sign change in the coefficients.
Proof of inequality (2.16): We set again $z=xy$, so that we have to show
$1+z^{p}+z^{2p}+y^{p}(1+z^{p})\geq(1+\frac{z}{y}+z)^{p}(1+y+z)^{p}\,.$
The restriction $x\leq y$ implies $0\leq z\leq y^{2}$. At the endpoints the inequality holds, by Lemma 2.9. As in the previous case, it is enough to check that the logarithmic derivative in $y$ of the ratio of left and right hand side vanishes at most once. It is
$\frac{p(1+z^{p})y^{p-1}}{1+z^{p}+z^{2p}+y^{p}(1+z^{p})}+\frac{pzy^{-1}}{z+(1+z)y}-\frac{p}{1+y+z}\,.$
Clearing all denominators, we obtain
$\quad y^{-1}\cdot pz(1+z)(1+z^{p}+z^{2p})$
$+y^{p-1}\cdot(p(1+z^{p})z(1+z)+p(1+z^{p})z(1+z))$
$+y^{p}\cdot(p(1+z^{p})(1+z)^{2}+p(1+z^{p})z+pz(1+z^{p})-pz(1+z^{p}))$
$+y\cdot(-p(1+z)(1+z^{p}+z^{2p}))\,.$
This is pseudo-polynomial in $y$ with sign sequence of its coefficients ++++++-+ + + -, so it has at most one zero. This completes the proof. ∎
Now we start the proof of Proposition 2.8.
Proof of Proposition 2.8.
The inequality (2.6) is homogeneous in $x$ and $y$, so we can assume that $x_{0}+x_{1}+x_{2}=y_{0}+y_{1}+y_{2}=1$. We denote
$\mathbf{P}=\{x\in[0,1]^{2}\,:\,\sum_{i=0}^{2}x_{i}=1\}\,.$
Lemma 2.11.
Suppose that $(x,y)\in\mathbf{P}^{2}$ minimizes
$J(x,y):=\sum_{k=0}^{4}\max_{i+j=k}x_{i}^{p}y_{j}^{p}$
on $\mathbf{P}^{2}$. Then for each $i$, we have $x_{i}=0$ or $x_{i}=1$ or there exists $k$ and $i^{\prime}$ such that $x_{i}y_{k-i}=x_{i^{\prime}}y_{k-i^{\prime}}$. Similarly, for each $j$ we have $y_{j}=0$, $y_{j}=1$ or there exists $k$ and $j^{\prime}$ such that $x_{k-j}y_{j}=x_{k-j^{\prime}}y_{j^{\prime}}$.
Proof of Lemma 2.11.
Suppose that $(x,y)$ is a minimizer which does not have the claimed property. By symmetry in $x$ and $y$, we can then assume that there exists $i$ such that $x_{i}\in(0,1)$ and such that $x_{i}y_{k-i}\neq x_{i^{\prime}}y_{k-i^{\prime}}$ for all $k=i,i+1,i+2$ and $i\neq i^{\prime}$. Now we perturb $x$ to decrease $J(x,y)$ as follows: We define
$x_{n}(t)=\begin{cases}(1+t)x_{n}&\text{if $n\neq i$}\\ x_{i}-t(1-x_{i})&\text{if $n=i$}\end{cases}\,.$
Since $x_{i}\in(0,1)$, there exists $\varepsilon>0$ such that $x(t)\in\mathbf{P}$ for all $t\in(-\varepsilon,\varepsilon)$. Furthermore, for each $k$ with $x_{i}y_{k-i}<\max_{i^{\prime}+j^{\prime}=k}x_{i^{\prime}}y_{j^{\prime}}$, there exists another $\varepsilon>0$ such that we have
$\max_{i^{\prime}+j^{\prime}=k}x_{i^{\prime}}(t)^{p}y_{j^{\prime}}=(1+t)^{p}\max_{i^{\prime}+j^{\prime}=k}x_{i^{\prime}}^{p}y_{j^{\prime}}$
for all $t\in(-\varepsilon,\varepsilon)$. Indeed, each product $x_{i^{\prime}}y_{j^{\prime}}$ with $i^{\prime}\neq i$ changes by the same factor, and the product $x_{i}y_{k-i}$ is strictly smaller. Finally, if $x_{i}y_{k-i}=\max_{i^{\prime}+j^{\prime}=k}x_{i^{\prime}}y_{j^{\prime}}$, then there also exists $\varepsilon>0$ such that we have
$x_{i}(t)^{p}y_{k-i}^{p}=\max_{i^{\prime}+j^{\prime}=k}x_{i^{\prime}}(t)^{p}y_{j^{\prime}}^{p}$
for all $t\in(-\varepsilon,\varepsilon)$, since $x_{i}y_{k-i}$ is stricly larger than all other products $x_{i^{\prime}}y_{k-i^{\prime}}$.
Taking the minimum of the $\varepsilon$’s, we obtain a neighbourhood $(-\varepsilon,\varepsilon)$ of 00 on which
$J(x(t),y)=\sum_{k=0}^{4}\max_{i^{\prime}+j^{\prime}=k}x_{i^{\prime}}(t)^{p}y_{j^{\prime}}^{p}$
is a concave function of $t$. Thus $(x,y)=(x(0),y)$ cannot be a minimizer. ∎
We now pick a minimizer $(x,y)$. We first assume that $x_{i},y_{j}\in(0,1)$ for all $i,j$. By Lemma 2.11, at least two equations $x_{i}y_{k-i}=x_{i^{\prime}}y_{k-i^{\prime}}$ must hold, involving each of $x_{0},x_{1}$ and $x_{2}$. It is easy to check that all such pairs of equations are
  1. 1)
    $x_{0}y_{1}=x_{1}y_{0}$ and $x_{0}y_{2}=x_{2}y_{0}$,
  2. 2)
    $x_{0}y_{1}=x_{1}y_{0}$ and $x_{1}y_{1}=x_{2}y_{0}$,
  3. 3)
    $x_{0}y_{1}=x_{1}y_{0}$ and $x_{1}y_{2}=x_{2}y_{1}$,
  4. 4)
    $x_{0}y_{2}=x_{1}y_{1}=x_{2}y_{0}$,
  5. 5)
    $x_{0}y_{2}=x_{1}y_{1}$ and $x_{1}y_{2}=x_{2}y_{1}$,
  6. 6)
    $x_{0}y_{2}=x_{2}y_{0}$ and $x_{1}y_{2}=x_{2}y_{1}$.
Pairs 1), 3) and 6) imply that $x=y$. In pair 2) there is no equation involving $y_{2}$. Adding any of the three possible equations one obtains that $x=y$. Pair 5) is symmetric to pair 2) after reversing the order of the variables, so it also reduces to $x=y$. Writing $x=y=(x_{0},x_{0}a,x_{0}ab)$ or $x=y=(x_{0},x_{0}b,x_{0}ba)$ with $0\leq a\leq b$, the inequality $J(x,y)\geq 1$ simplifies to (2.14) and (2.15), respectively, so it holds by Lemma 2.10.
The only remaining case is pair 4). In that case we can write $x=(x_{0},ax_{0},abx_{0})$ and $y=(y_{0},by_{0},aby_{0})$, and by symmetry we have without loss of generality $0\leq a\leq b$. Then the inequality simplifies to (2.16), so follows again from Lemma 2.10.
Now we consider minimizers $(x,y)$ on the boundary of $\mathbf{P}^{2}$. If $(x,y)$ is on the boundary, then at least one of the variables is 00. Since $x,y$ are nondecreasing, this must be $x_{0}$ or $y_{0}$. We assume without loss of generality that $y_{0}=0$.
If $y_{1}=0$ or then $y=(0,0,1)$, and the inequality holds. So we can assume that $y_{1},y_{2}\in(0,1)$. If any of $x_{0},x_{1},x_{2}$ equals $1$, the inequality is trivial. If $x_{0}=0$, then the inequality reduces to the case $m=1$, which is proven in (2.3). If $x_{0}\neq 0$ then $x_{0},x_{1},x_{2}\in(0,1)$ since $x$ is nondecreasing. So by Lemma 2.11, one of the $6$ pairs of equations listed above must hold. Pairs 1), 2), 3), 4) and 6) contradict the assumption that $y_{0}=0$ and all other variables are not 00. If pair 5) holds, then we can write
$x=(x_{0},ax_{0},a^{2}x_{0})\,,\quad\quad y=(0,y_{1},ay_{1})\,.$
Then the inequality reduces to (2.10) above, so it follows from Lemma 2.9. ∎
2.3. The proof of Theorem 1.4
By Remark 3 and Lemma 11.4 in [13] (which is similar to Lemma 2.2) it suffices to show that for all non-negative finitely supported $g,h:\mathbb{Z}\to[0,\infty)$ and $f(x)=1_{\{0\}}+\delta 1_{\{1\}}$ with any $\delta\in[0,1]$, we have
$\frac{\|f\bar{*}g\bar{*}h\|_{1}}{\|g\|_{p}\|h\|_{q}}\geqslant(1+\delta^{c})^{1/c}.$
Theorem 11.1 in [13] implies that
$\frac{\|f\bar{*}g\bar{*}h\|_{1}}{\|g\|_{p}\|h\|_{q}}\geqslant c_{\delta}:=\frac{(1-\delta^{p})^{1/p}(1-\delta^{q})^{1/q}}{1-\delta}.$
The case $\delta=1$ should be understood as $\lim_{\delta\to 1}c_{\delta}=p^{1/p}q^{1/q}$. Thus it suffices prove the following subtle inequality.
Lemma 2.12.
For any $p,q>1$ satisfying $11/p+1/q=$ 1 one has
$\frac{(1-x^{p})^{1/p}(1-x^{q})^{1/q}}{(1-x)(1+x^{c})^{1/c}}\geqslant 1,$
for any $x\in[0,1]$, where $c:=1/\log_{2}{(p^{1/p}q^{1/q})}$ is chosen in such a way that $x\rightarrow 1$ gives an equality.
Proof.
Denote the left hand side of the inequality by $F(x)$. We want to prove that $\log{F(x)}\geqslant 0$ where $\log{F(x)}$ is given by
$\log{F(x)}=\tfrac{1}{p}\log{|1-x^{p}|}+\tfrac{1}{q}\log{|1-x^{q}|}-\log{|1-x|}-\tfrac{1}{c}\log{|1+x^{c}|}.$
We have
$x\cdot\partial_{x}\log{F(x)}$ $=-\frac{x^{p}}{1-x^{p}}-\frac{x^{q}}{1-x^{q}}+\frac{x}{1-x}-\frac{x^{c}}{1+x^{c}}$
$=\frac{1}{1-x}+\frac{1}{1+x^{c}}-\frac{1}{1-x^{p}}-\frac{1}{1-x^{q}},$
and so $H(x):=x(1-x)(1+x^{c})(1-x^{p})(1-x^{q})\cdot\partial_{x}\log{F(x)}$ can be written as a pseudo-polynomial
$x-x^{c}+2x^{1+c}-x^{p}-x^{q}-x^{1+c+p}-x^{1+c+q}+2x^{p+q}-x^{p+q+1}+x^{p+q+c}.$
We have $H(1/x)=H(x)\cdot x^{-1-c-p-q}$ and by the rule of signs $H(x)$ has at most 6 positive roots. Indeed, it’s easy to see that $c\geqslant 1$ by Jensen’s inequality, and so the smallest exponent is $1$ and the largest is $x^{p+q+c}$. When arranged in the increasing order of exponents, they have at most one neighbouring term with a negative sign each, and the remaining two terms with a plus sign (i.e. $2x^{1+c}$ and $2x^{p+q}$) give at most 4 more changes of sign.
Note that as $x\rightarrow 0$ we have
$F(x)=\frac{(1-x^{p}/p)(1-x^{q}/q)}{(1-x)}\cdot(1+o_{x}(1))=\frac{1-o(x)}{1-x},$
and so $F(x)\geqslant 1$ in some neighbourhood of zero.
Now assume that $F(t)<1$ for some $t\in(0,1)$. Then $H(t)$ must have a root on both $(0,t)$ and $(t,1)$. Hence, together with a $x\mapsto 1/x$ symmetry and the triple root at zero we have at least 7 positive roots leading to a contradiction. ∎
2.4. The proof of Theorem 1.5
The following lemma plays an important role in the proof of Theorem 1.5, it is motivated by Theorem 1.1 in [5] and can be established with a similar strategy.
Lemma 2.13.
Let $d\geqslant 1$, $k\in\mathbb{N}$, $A_{1},A_{2},\dots,A_{k}\subseteq\mathbb{Z}^{d}$ be finite sets, and suppose that $U=\{0,1,\dots,k-1\}^{d}$. Then $|A_{1}+A_{2}+\dots+A_{k}+U|\geqslant|U|\prod_{i=1}^{k}|A_{i}|^{1/k}$.
Notice that the inequality in Lemma (2.13) is optimal, in fact if $A_{1}=A_{2}=\dots=A_{k}=U=\{0,1,\dots,k-1\}^{d}$, then, both sides are equal to $k^{2d}$. Before we verify the lemma let us complete the proof of Theorem 1.5.
The upper bound in Theorem 1.5 can be obtained by considering the case $A=\{0,1,\dots,k-1\}^{d}$. We will focus our attention on the lower bound. Let $\alpha:=\frac{|kA|}{|A|}$. By the Plünnecke inequality for different summands in commutative group [4, Theorem 1.1] there is a non-empty subset $B$ of $A$ such that
$\alpha^{\frac{k}{k-1}}|B|\geqslant|B+kA|.$
Then, by Lemma 2.13 we have
$\alpha^{\frac{k}{k-1}}|B|$ $\geqslant|B+kA|$
$\geqslant|B+(k-1)A+\{0,1,\dots,k\}^{d}|$
$\geqslant k^{d}|B|^{\frac{1}{k}}|A|^{\frac{k-1}{k}}.$
Therefore
$\alpha^{\frac{k}{k-1}}|B|^{\frac{k-1}{k}}\geqslant k^{d}|A|^{\frac{k-1}{k}}\geqslant k^{d}|B|^{\frac{k-1}{k}}.$
Thus, we conclude that
$\alpha\geqslant k^{\frac{(k-1)d}{k}}.$
2.4.1. Proof of Lemma 2.13
The proof of this lemma follows from an adaptation of the proof of [5, Theorem 1.2] by Green, Matolcsi, Rusza, Shakan and Zhelezov. Similarly to [5], we obtain Lemma 2.13 by discretizing the continuous Prékopa–Leindler inequality for many functions.
Proposition 2.14.
Let $k\in\mathbb{N}$, $a_{1},a_{2},\dots,a_{k}:\mathbb{Z}\rightarrow[0,\infty)$ be compactly supported functions, let $\lambda>0$ and $p_{j}:=e^{j\lambda}\frac{e^{\lambda}-1}{e^{k\lambda}-1}$ for all $j\in\{0,1,\dots,k-1\}$. Then we have
$\sum_{n\in\mathbb{Z}}\max_{j\in\{0,1,\dots,k-1\}}(p_{j}a_{1}\overline{\ast}a_{2}\overline{\ast}\dots\overline{\ast}a_{k}(n-j))\geqslant\prod_{i=1}^{k}\|a_{i}\|_{k}.$ (2.18)
Corollary 4.
In particular, taking the limit in the left hand side of (2.18) as $\lambda\to 0^{+}$, we obtain that (2.18) holds when $p_{0}=p_{1}=\dots=p_{k-1}=\frac{1}{k}$.
Assume that the functions $f_{i}:\mathbb{R}\rightarrow[0,\infty)$ are compactly supported and piece-wise continuous for each $i\in\{1,2,\dots,k\}$. Then the (1-dimensional) generalized Prékopa–Leindler inequality states that
$\int f_{1}\overline{\ast}f_{2}\overline{\ast}\dots\overline{\ast}f_{k}\geqslant k\prod_{i=1}^{k}\|f_{i}\|_{k},$ (2.19)
where the left hand side denotes the max-convolution, which is defined by
$f_{1}\overline{\ast}f_{2}\overline{\ast}\dots\overline{\ast}f_{k}(x):=\sup_{x_{1}+x_{2}+\dots+x_{k};x_{i}\in\mathbb{R}}\prod_{i=1}^{k}f_{i}(x_{i}),$
and as usual
[…]
Proof of Proposition 2.14.
We use (2.19) with the functions $f_{i}$ defined by
$f_{i}(x):=e^{\lambda\{x\}}a_{i}(\lfloor x\rfloor),$
for all $i\in\{1,2,\dots,k\}$.
Let $n\in\mathbb{Z}$ and $10\leqslant t<$ 1. Assume that $\sum_{i=1}^{k}x_{i}=n+t$. Then, since $x-1<\lfloor x\rfloor\leqslant x$, we have $n-k<\sum_{i=1}^{k}\lfloor x_{i}\rfloor<n+1$. If $\sum_{i=1}^{k}\lfloor x_{i}\rfloor=n-m$ for some $m\in\{0,1,\dots,k-1\}$, then
$\prod_{i=1}^{k}f_{i}(x_{i})\leqslant e^{\lambda(t+m)}a_{1}\overline{\ast}a_{2}\overline{\ast}\dots\overline{\ast}a_{k}(n-m).$
Thus
$f_{1}\overline{\ast}f_{2}\overline{\ast}\dots\overline{\ast}f_{k}(n+t)\leqslant e^{\lambda t}\max_{m\in\{0,1,\dots,k-1\}}e^{\lambda m}a_{1}\overline{\ast}a_{2}\overline{\ast}\dots\overline{\ast}a_{k}(n-m).$
Taking integrals from 0 to 1 with respect to $t$ and then adding over all $n\in\mathbb{Z}$ we obtain
[…]
On the other hand,
$\|f_{i}\|_{k}^{k}=\frac{e^{k\lambda}-1}{k\lambda}\|a_{i}\|_{k}^{k},$
for each $i\in\{1,2,\dots,k\}$. Therefore, from (2.19) we obtain
$\sum_{n\in\mathbb{Z}}\max_{m\in\{0,1,\dots,k-1\}}e^{\lambda m}a_{1}\overline{\ast}a_{2}\overline{\ast}\dots\overline{\ast}a_{k}(n-m)\geqslant\frac{e^{k\lambda}-1}{e^{\lambda}-1}\prod_{i=1}^{k}\|a_{i}\|_{k}.$
The result then follows from the fact that $p_{j}:=e^{j\lambda}\frac{e^{\lambda}-1}{e^{k\lambda}-1}$. ∎
Proof of Lemma 2.13.
We proceed by induction on $d$. The proof of the inductive step also proves the base case $d=1$.
Recall that $U=\{0,1,\dots,k-1\}^{d}$. We define $\pi:U\to\{0,1,\dots,k-1\}$ to be the projection onto the last coordinate, i.e $\pi(x_{1},\dots,x_{d})=x_{d}$ for all $(x_{1},\dots,x_{d})\in\{0,1,\dots,k-1\}^{d}$. For each $i\in\{1,2,\dots,k\}$ and $x\in\{0,1,\dots,k-1\}$, let $A_{i,x}:=A_{i}\cap\pi^{-1}(x)$ be the fibre of $A_{i}$ above $x$. The set $U$ has $k$ fibres $U_{0},U_{1},\dots,U_{k-1}$ and they are all $d-1$-dimension cubes, here $U_{i}=\{0,1,\dots,k-1\}^{d-1}\times\{i\}$ for each $i\in\{0,1,\dots,k-1\}$. Moreover, we have that $\frac{|U_{i}|}{|U|}=\frac{1}{k}$ for each $i\in\{0,1,\dots,k-1\}$.
We note that the fibre of $A_{1}+A_{2}+\dots+A_{k}+U$ above $n$ contains the set $A_{1,x_{1}}+A_{2,x_{2}}+\dots+A_{k,x_{k}}+U_{m}$ whenever $x_{1}+x_{2}+\dots+x_{k}=n-m$ for $m\in\{0,1,\dots,k-1\}$. By induction,
$|A_{1,x_{1}}+A_{2,x_{2}}+\dots+A_{k,x_{k}}+U_{m}|\geqslant|U_{m}|\prod_{i=1}^{k}|A_{i,x_{i}}|^{1/k},$
for all $m\in\{0,1,\dots,k-1\}$.
For each $i\in\{1,2,\dots,k\}$ we define $a_{i}:\mathbb{Z}\to[0,+\infty)$ by $a_{i}(x):=|A_{i,x}|^{1/k}$ for all $x\in\{0,1,\dots,k-1\}$ and $a_{i}(x)=0$ otherwise. Then summing over $n$, using the fact that $|U_{m}|=\frac{|U|}{k}$ for all $10\leqslant m\leqslant k-$ 1, and using Proposition 2.14, we obtain
$|A_{1}+A_{2}+\dots+A_{k}+U|$
$=\sum_{n\in\mathbb{Z}}|(A_{1}+A_{2}+\dots+A_{k}+U)_{n}|$
$\geqslant\sum_{n}\frac{|U|}{k}\max_{x_{1}+x_{2}+\dots+x_{k}=n-m}\prod_{i=1}^{k}|A_{i,x_{i}}|^{1/k}$
$\geqslant|U|\sum_{n}\max_{m\in\{0,1,\dots,k-1\}}\big{(}\frac{1}{k}a_{1}\overline{\ast}a_{2}\overline{\ast}\dots\overline{\ast}a_{k}(n-m)\big{)}$
$\geqslant|U|\prod_{i=1}^{k}\|a_{i}\|_{k}=|U|\prod_{i=1}^{k}|A_{i}|^{1/k}.$
3. Acknowledgments
The authors are very thankful to Jaume de Dios Pont, Pavlos Kalantzopoulos, Fedor Petrov, and Terence Tao for helpful discussions. L.B. was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813 as well as SFB 1060. P.I. acknowledges support from NSF grant CAREER-DMS-2152401.

### v2

@@
 Abstract.
-We show that for all subsets $A_{1},A_{2},\dots,A_{n}\subseteq\{0,1\}^{d}$, we have
-$|A_{1}+A_{2}+\dots+A_{n}|\geqslant\prod_{i=1}^{n}|A_{i}|^{(\log_{2}(n+1))/n}.$
-In addition, we show that for all $A,B\subseteq\{0,1,2\}^{d}$ we have
+We show that for all $A,B\subseteq\{0,1,2\}^{d}$ we have
 $|A+B|\geqslant(|A||B|)^{\log(5)/(2\log(3))}.$
@@
 holds for all $p\in(1,\infty)$, where $q=\frac{p}{p-1}$ is the conjugate exponent of $p$.
-All three estimates are dimension free with the best possible exponents.
-We also discuss applications to various related problems.
+All the estimates are dimension free with the best possible exponents.
+We discuss applications to various related problems.
 2020 Mathematics Subject Classification:
@@
 $\frac{\log(nm+1)}{n\log(m+1)}\geqslant p_{n,m}\,.$ (1.5)
-3. The main result in [17] states that $p_{2,1}=\frac{\log(3)}{\log(4)}$, i.e., we have equality in (1.5) in the case $(n,m)=(2,1)$.
+3.
+The main result in [21] and independently in [9] states that $p_{2,1}=\frac{\log(3)}{\log(4)}$, i.e., we have equality in (1.5) in the case $(n,m)=(2,1)$.
+The equality $p_{n,1}=\frac{\log(n+1)}{n\log(2)}$ was conjectured in [9], and established [^1] We are grateful to Ben Green for pointing to us Appendix B in [6] which contains more comprehensive list of references on the value $p_{n,1}$.
+in [2] and independently in [16].
+The proof in [16] uses conformal maps and Hadamard’s three circle theorem, whereas the proof in [2] is more direct and uses a clever inductive argument.
+We will present a new proof (Theorem 1.4) of the equality $p_{n,1}=\frac{\log(n+1)}{n\log(2)}$ based on Descartes’ rule of signs.
 4. For $m\geqslant 3$ let $\tau_{m}$ be the solution of the equation
@@
 $f\bar{*}g(z)=\sup_{x\in\mathbb{Z}^{d}}f(x)g(z-x).$
-In the next theorem we show that we have equality in (1.5) for the pairs $(n,m)=(n,1)$ for all $n\geqslant 1$.
-Theorem 1.1.
-For all $d\geqslant 1$, and all $f_{j}:\mathbb{Z}^{d}\to[0,\infty)$, $j=1,\ldots,n$, supported on $\{0,1\}^{d}$, we have
-[…]
-In particular, for all subsets $A_{j}\subseteq\{0,1\}^{d}$, we have
-$\left|\sum_{j=1}^{n}A_{j}\right|\geqslant\left(\prod_{j=1}^{n}|A_{j}|\right)^{p}$ (1.7)
-with the best possible power $p=\frac{\log_{2}(n+1)}{n}$.
-The inequality (1.7) follows from (1.6) by taking $f_{j}=1_{A_{j}}$ for each $j=1,\ldots,n$. A related inequality to (1.6) was established by Ivanisvili [8, Theorem 1].
-Next we affirmatively answer the question asked in [11] which also confirms the case $m=2$ of the question in [1].
+In the next theorem we affirmatively answer the question asked in [14] which also confirms the case $m=2$ of the question in [1].
 Theorem 1.1.
 For all $d\geqslant 1$, and all $f,g:\mathbb{Z}^{d}\to[0,\infty)$ supported on $\{0,1,2\}^{d}$, we have
-[…]
+$\|f\bar{*}g\|_{1}\geqslant\|f\|_{q}\|g\|_{q},\quad\text{where}\quad q=\frac{2\log(3)}{\log(5)}.$ (1.6)
 In particular, for all $A,B\subset\{0,1,2\}^{d}$, and all $d\geqslant 1$ we have
-$|A+B|\geqslant|A|^{p}|B|^{p}$
+$|A+B|\geqslant|A|^{p}|B|^{p}$ (1.7)
 holds with the best possible power $p=\frac{\log(5)}{2\log(3)}=0.732\ldots$.
+The inequality (1.7) follows from (1.6) by taking indicator functions of sets i.e., $f=1_{A}$ and $g=1_{B}$.
 The following result was obtained by Matolcsi, Ruzsa, Shakan and Zhelezov.
@@
 A short self-contained proof of Theorem 1.2 was recently given in [7] in the case $p=q=2$.
+The next theorem due to Landau, Logan, Shepp [16] and independently Brown, Keane, Moran and Pierce [2] states that we have equality in (1.5) for the pairs $(n,m)=(n,1)$ for all $n\geqslant 1$.
+We will present a new proof of it based on Decartes’ rule of signs.
+Theorem 1.4 ([16], [2]).
+For all $d\geqslant 1$, and all $f_{j}:\mathbb{Z}^{d}\to[0,\infty)$, $j=1,\ldots,n$, supported on $\{0,1\}^{d}$, we have
+$\|f_{1}\bar{*}\cdots\bar{*}f_{n}\|_{1}\geqslant\prod_{j=1}^{n}\|f_{j}\|_{q},\quad\text{where}\quad q=\frac{n}{\log_{2}(n+1)}.$ (1.12)
+In particular, for all subsets $A_{j}\subseteq\{0,1\}^{d}$, we have
+$\left|\sum_{j=1}^{n}A_{j}\right|\geqslant\left(\prod_{j=1}^{n}|A_{j}|\right)^{p}$ (1.13)
+with the best possible power $p=\frac{\log_{2}(n+1)}{n}$.
+A related inequality to (1.12) was established by Ivanisvili [11, Theorem 1].
 For any positive integer $k$ set
@@
 Remark 2.
-Our Theorem 1.5 can be seen as an extension of a result obtained by Green and Tao [6, Proposition 4.1], for $k=2$.
+Our Theorem 1.5 can be seen as an extension of a result obtained by Green and Tao [8, Proposition 4.1], for $k=2$.
 For any $\lambda>0$ set
@@
 The additive energy $E(A)$ of a finite subset $A$ of an additive group $G$ is defined as $E(A):=|\{(a_{1},a_{2},a_{3},a_{4})\in A^{4};\ a_{1}+a_{2}=a_{3}+a_{4}\}|,$ (see [20]).
-It was proved by Kane and Tao [10, Theorem 7] that $E(A)\leqslant|A|^{\log_{2}6}$ for all $A\subseteq\{0,1\}^{d}$.
-For any $k\geqslant 2$, the $k-$additive energy $E_{k}(A)$ of a finite subset $A$ of an additive group $G$ is defined as $E_{k}(A):=|\{(a_{1},a_{2},\dots,a_{2k})\in A^{2k};\ a_{1}+a_{2}+\dots+a_{k}=a_{k+1}+a_{k+2}+\dots+a_{2k}\}|.$ The result of Kane-Tao was later extended to all $k\geqslant 2$, see [2, Theorem 2] and [12, Corollary 2], establishing that $E_{k}(A)\leqslant|A|^{\log_{2}{{2k}\choose{k}}}$ for all $A\subseteq\{0,1\}^{d}$.
-Combining this estimate with our Theorem 1.1 we obtain
+It was proved by Kane and Tao [13, Theorem 7] that $E(A)\leqslant|A|^{\log_{2}6}$ for all $A\subseteq\{0,1\}^{d}$.
+For any $k\geqslant 2$, the $k-$additive energy $E_{k}(A)$ of a finite subset $A$ of an additive group $G$ is defined as $E_{k}(A):=|\{(a_{1},a_{2},\dots,a_{2k})\in A^{2k};\ a_{1}+a_{2}+\dots+a_{k}=a_{k+1}+a_{k+2}+\dots+a_{2k}\}|.$ The result of Kane-Tao was later extended to all $k\geqslant 2$, see [3, Theorem 2] and [15, Corollary 2], establishing that $E_{k}(A)\leqslant|A|^{\log_{2}{{2k}\choose{k}}}$ for all $A\subseteq\{0,1\}^{d}$.
+Combining this estimate with (1.13) we obtain
 $E_{k}(A)$ $\leqslant|A|^{\log_{2}{{2k}\choose{k}}}=|A|^{(\log_{2}{\frac{k+1}{2}})(\log_{\frac{k+1}{2}}{{2k}\choose{k}})}$
@@
 In [10] Hanson and Petridis study sum of dilates, they proved that if $A$ is a subset of a commutative group $G$ and $K$ is a parameter such that $|A+A|\leqslant K|A|$, then $|A+2\cdot A|\leqslant K^{2.95}|A|$.
-As a consequence of the simple inequality $\frac{|A+2\cdot A|}{|A|}\leqslant|A|$ and our Theorem 1.1 we obtain
+As a consequence of the simple inequality $\frac{|A+2\cdot A|}{|A|}\leqslant|A|$ and (1.13) one has
 Corollary 2.
@@
 As in [4] the hypercube $\{0,1\}^{n}\subset\mathbb{R}^{n}$ can be projected (which is also bijection onto its image) by a transformation $T$ onto the real line $\mathbb{R}$ giving a set $P_{n}$ of cardinality $N=2^{n}$ so that any subset of $\tilde{A}\subset P_{n}$ has the same number of distinct m-sums as the set $A\subset\{0,1\}^{n}$, where $T(A)=\tilde{A}$.
-Thus the corollary is a direct consequence of Theorem 1.1.
+Thus the corollary is a direct consequence of (1.13).
 The projection is defined as follows: Let $n\in\mathbb{Z}^{+}$.
@@
 Let $\alpha:=\frac{|kA|}{|A|}$.
-By the Plünnecke inequality for different summands in commutative group [4, Theorem 1.1] there is a non-empty subset $B$ of $A$ such that
+By the Plünnecke inequality for different summands in commutative group [5, Theorem 1.1] there is a non-empty subset $B$ of $A$ such that
 $\alpha^{\frac{k}{k-1}}|B|\geqslant|B+kA|.$
@@
 2.4.1. Proof of Lemma 2.13
-The proof of this lemma follows from an adaptation of the proof of [5, Theorem 1.2] by Green, Matolcsi, Rusza, Shakan and Zhelezov.
+The proof of this lemma follows from an adaptation of the proof of [7, Theorem 1.2] by Green, Matolcsi, Rusza, Shakan and Zhelezov.
 Similarly to [7], we obtain Lemma 2.13 by discretizing the continuous Prékopa–Leindler inequality for many functions.
@@
 3. Acknowledgments
-The authors are very thankful to Jaume de Dios Pont, Pavlos Kalantzopoulos, Fedor Petrov, and Terence Tao for helpful discussions.
+The authors are very thankful to Ben Green, Jaume de Dios Pont, Pavlos Kalantzopoulos, Fedor Petrov, and Terence Tao for helpful discussions.
 L.B.